OpenStudy (anonymous):

Estimate the circumference. Use 3.14 or to approximate π.

OpenStudy (anonymous):



OpenStudy (anonymous):

will fan and medal

jhonyy9 (jhonyy9):

so 40 is the radius ?

OpenStudy (anonymous):

no 140

OpenStudy (anonymous):

hello??

jhonyy9 (jhonyy9):

do you know formula for circumference of a circle ?

OpenStudy (anonymous):

not really

jhonyy9 (jhonyy9):

ok

jhonyy9 (jhonyy9):

formula you know ?

jhonyy9 (jhonyy9):

so this is C = 2pi*r

OpenStudy (anonymous):

k so

jhonyy9 (jhonyy9):

now just substitute the knowed value of pi and r and make the calcul

OpenStudy (anonymous):

how do you do that?

jhonyy9 (jhonyy9):

C= 2*3,14 *140 = ? ok ?

jhonyy9 (jhonyy9):

just multiplie 2*3,14 =6,28

jhonyy9 (jhonyy9):

and 6,28*140 = ?

OpenStudy (anonymous):

is it 6.28 times 140 right which equals 879.2 right???

jhonyy9 (jhonyy9):

exactly nice so this is the length of circumference of this circle was easy ,yes ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

tks

jhonyy9 (jhonyy9):

so good luck bye

OpenStudy (anonymous):

wait how can i fan you??